Jan 31, 2012 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST'

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Monitoring
    • Send additional SiPMs to Athens for tests (Elton). Electronics is out for fabrication
  2. TDR
    • Create outline to assign writing homework (Elton)

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Collaboration Meeting
    1. Friday at 2:30pm, 4 30 minute talks.
  4. Monitoring.
  5. FCAL
    1. Construction Status: mechanical, bases (notes from Matt)
    2. Beam tests in Hall B (John)
  6. BCAL
    1. BCAL Beam Tests 2012
  7. USM update
    1. Status report
    2. Status Stage 3 Analysis
  8. TDR update
  9. Simulations/Reconstruction
  10. Any other business


Attending: David, Tim, Yi, John, Beni (JLab), Matt, Kei (IU), Serguei, Hayk, Javiera (USM), George, Andrei, Zisis (Regina), Regrets: Athens"

  1. Announcements: none
  2. Review of Action Items: Elton was not present so we did not get an update on the items. Once Elton provides the outline for the TDR, we agreed that persons should be assigned to each section and we should aim for having a draft by the June Collaboration Meeting.
  3. Collaboration Meeting
    1. FCAL Construction (Matt) - 30min
    2. FCAL Beam Tests (John) - 30min
    3. BCAL Beam Tests (Zisis) - 15min
    4. SiPM (?) - 45 min (the JLab, USM, Regina work and results should be organized into a coherent presentation; will work this out next week during the readout meeting).
    5. BCAL Reconstruction Code: Matt and Will L. will work offline to prepare this; if short, it will be shown by Zisis; otherwise moved to the Offline Session.
  4. Monitoring: No report, but things are proceeding nominally.
  5. FCAL
    1. Construction Status:
      1. Aluminum insert is in the shop; welded steel frame provided by Tim and was sent to a shop for a quote (expected in a week) before going to bid process; magnetic shield order placed and expecting a March delivery
      2. Light guide gluing on PMTs around 1000 done; 900 tested, 100 more need checks. Old PMT photocathode window is cracking after LG is glued; thermal cycle is not the reason (tested), maybe handling?
      3. Cockroft-Walton bases: working on HV multiplier board; still issues with conformal coating, hoping that next batch would have solved most: process seems to be converging.
    2. Beam tests in Hall B (John): gain balance is now believable, 22% energy resolution at ~100 MeV, overall going well
  6. BCAL
    1. BCAL Beam Tests 2012: Tim said that 80 SiPM units should be available by end of February. John said that HallB has e- running until end of Feb with a short installation period afterwards. Tim mentioned that the updated schedule will be available later today. Sascha is planning a full, separate DAQ for the BCAL (unless FCAL is done by that point) and there should be plenty of flash250 boards available. The default mode of operation should be planned to be "standalone" with our own trigger counters. John mentioned that "characterized" scintillators will be worked out at three positions for the FCAl and would be useable by the BCAL. FCAL is still in the alcove and will move to the floor during the install time.
  7. USM update
    1. Status report Javiera presented pictures of the stages (stations) and information including serial no. labelling/tracking. The microscope/camera (stage 1) works with software to measure the active area of each MPPC. Very nice thermocouple results were shown, with very small differences between the 16 cell sites.
    2. Status Stage 3 Analysis Hayk showed improved plots where some issues associated with the gain were resolved since last week. There are still some issues that need remeasuring or looking into more detail (e.g gain of specific cells, cross talk, QDC pedestal). The work is progressing well. One concern is that these measurements are with low number of Npe. We should discuss higher Npe response, more in line with what the experiment will see over most regions of acceptance.
  8. TDR update
  9. Simulations/Reconstruction: no report.
  10. Any other business: none